OpenStudy (liv1234):

let f(x)=-5x-4 and g(x)=6x-7. find f(x)+g(x) A. -11x+3 B. x+3 C. -11x-11 D. x-11 My answer is D, can someone check to make sure I'm correct please?

OpenStudy (kuoministers):

Yepp All good

OpenStudy (liv1234):

Can you check another one for me?

OpenStudy (kuoministers):

yeah ok

OpenStudy (liv1234):

2. let f(x)=3x+2 and g(x)=6x-7. find f(x)-g(x) A. 3x-9 B. -3x+9 C. -3x-5 D. 2x-1 My answer is C.

OpenStudy (kuoministers):

Should be B

OpenStudy (kuoministers):

2 - (-7) = 9

OpenStudy (liv1234):

Ohh, I was just about to correct myself! Do you mind checking one more for me?

OpenStudy (kuoministers):

yeah sure last one

OpenStudy (liv1234):

3. let f(x)=-5+3 and g(x)=6x-2. find f*g and its domain A. -30x^2+28x-6; all real numbers except x=1/3 B. -15x^2-8x-12; all real numbers except x=3/5 C. -30x^2+28x-6; all real numbers D. -15x^2-8x-12; all real numbers My answer is A.

OpenStudy (kuoministers):

do you mean -5x+3 = f(x) ?

OpenStudy (liv1234):

Yes.

OpenStudy (liv1234):

f(x) = -5x+3

OpenStudy (kuoministers):

Thats good, but the domain is X all real

OpenStudy (kuoministers):

Other than that good work! :D

OpenStudy (liv1234):

So, it would be A?

OpenStudy (kuoministers):

Would be C

OpenStudy (liv1234):

Ohh, okay, thank you!

OpenStudy (kuoministers):

Also... sub x = 1/3 if you wanted to check. You can get an ans

OpenStudy (kuoministers):

No problem. Happy to help

OpenStudy (liv1234):

Thank you very much!(: